# Closed-Form Feedback-Free Learning with Forward Projection

### v2

@@ Abstract
State-of-the-art methods for backpropagation-free learning employ local error feedback to direct iterative optimisation via gradient descent. In this study, we examine the more restrictive setting where retrograde communication from neuronal outputs is unavailable for pre-synaptic weight optimisation. To address this challenge, we propose Forward Projection (FP). This novel randomised closed-form training method requires only a single forward pass over the entire dataset for model fitting, without retrograde communication. Our key innovation is generating target values for pre-activation membrane potentials at each layer via nonlinear projections of pre-synaptic inputs and the labels. Our approach generates target potentials from randomised encodings combining information from inputs and labels. Local loss functions are optimised over pre-synaptic inputs using closed-form regression, without feedback from neuronal outputs or downstream layers. Interpretability is a key advantage of FP training; membrane potentials of hidden neurons in FP-trained networks encode information which are interpretable layer-wise as label predictions. We demonstrate the effectiveness of FP across four biomedical datasets, comparing it with backpropagation and local learning techniques such as Forward-Forward training and Local Supervision in multi-layer perceptron and convolutional architectures. In some few-shot learning tasks, FP yielded more generalisable models than those optimised via backpropagation. In large-sample tasks, FP-based models achieve generalisation comparable to gradient descent-based local learning methods while requiring only a single forward propagation step, achieving significant speed up for training. Interpretation functions defined on local neuronal activity in FP-based models successfully identified clinically salient features for diagnosis in two distinct datasets – identification of key signatures of myocardial infarction in time-series data from electrocardiogram sequences and detection of regions of choroid neovascularisation in optical coherence tomography images.
@@ Results
Forward Projection was compared to other local learning methods, including random features (RF) [23], Local Supervision (LS) [9], and Forward-Forward (FF) [15] in multiple tasks using equivalent model architectures (Table 1). The Fashion MNIST (FMNIST) image classification task was modelled on a multi-layer perceptron (MLP). Forward projection achieved higher test accuracy than other local learning methods in this task, approaching the performance of the backpropagation reference standard. Two large-scale biomedical sequence modelling tasks were also evaluated with one-dimensional convolutional neural network (1d-CNN) architectures. The Promoters task required the identification of human non-TATA promoters, a class of gene promoter regions which increase transcription of DNA sequences [26, 27]. Forward projection yielded higher test performance than all other local learning methods in this task. The PTBXL-MI task [28] required diagnosis of myocardial infarction (MI), a heart condition commonly known as “heart attack”, from 12-channel electrocardiograph (ECG) recordings. FP and LS performed comparably in this task. FP also outperformed all other methods in the optimisation of models with more complex neuronal activations such as modulo and polynomial activation functions (see remark A.11 in SI). FP was implemented in a transformer architecture to discriminate the first two classes of the CIFAR10 dataset, significantly outperforming random features. Backpropagation performance benchmarks in PTBXL-MI and Promoters tasks were comparable to previous studies [29, 27].
@@
Table 1: Test performance of various learning methods across different datasets. The FMNIST dataset was modelled with a multi-layer perception; Promoters and PTBXL-MI datasets were modelled using 1D convolutional neural networks (1D-CNNs); and CIFAR was modelled using vision transformers. BP: bacpropagation; DTP: Difference Target Propagation; $F$: Forward Forward; FP: Forward Projection; LS: Local Supervision; PC: Predictive Coding; RF: random features
@@
Figure 2: A: Comparison of feedback-free fitting methods on FMNIST. MLP architectures had 1000 neurons in the first and second layers and 100, 200, 400 or 800 neurons in the final hidden layer. B-D: Test performance of few-shot trained 2D-CNN models. Mean test AUC is reported over 50 few-shot training experiments. B: Chest X-ray (CXR) task. C: Optical Coherence Tomography (OCT) task. D: CIFAR task in which models were required to classify the first two classes (aeroplane and automobile). Mean test AUC is reported over 50 few-shot training experiments. Models were fitted with $N\in\{5,10,15,20,30,40,50\}$ training samples from each class in CXR and OCT tasks and $N\in\{25,50,75,100\}$ samples per class for the CIFAR task. Predictive Coding and Difference Target Propagation are plotted separately in Supplementary Figure 7. BP: backpropagation; $F$: Forward-Forward; FP: Forward Projection; LS: Local Supervision; RF: random features.
@@
“Few-shot” learning is a constrained learning scenario in which the number of data samples available for training is small. High-dimensional data, such as images, pose a challenge for few-shot learning, as many spurious features may exist. Thus, successful few-shot training methods must select generalisable features in the presence of these confounders. Few-shot learning was assessed in image classification tasks using a 2D-CNN architecture. The optical coherence tomography task (OCT) [30] required that the model discriminate images of healthy retinas from those with choroid neovascularization, a pathology that affects the eye and manifests mainly as abnormal growth of blood vessels behind the retina. The paediatric chest X-ray (CXR) task [30] required the model to discriminate between images of viral pneumonia, bacterial pneumonia, or healthy controls. The CIFAR task required the model to discriminate between the first two CIFAR10 classes (aeroplane and automobile). Few-shot training datasets were generated by subsampling, with $N\in\{5,10,15,20,30,40,50\}$ training examples from each class for OCT and CXR tasks, and $N\in\{25,50,75,100\}$ training examples per class for the CIFAR task. Data augmentation was not employed. Model generalisability was assessed on all test data (CXR: $N_{\text{test}}=431$; OCT: $N_{\text{test}}=327$, CIFAR: $N_{\text{test}}=2000$). Forward Projection-trained models demonstrated the greatest few-shot generalisability in CXR (Figure 2-B), OCT (Figure 2-C), outperforming all other methods, including backpropagation, in all the tasks with $N\leq 40$ training samples. In the CIFAR few-shot task, Forward Projection achieved the best performance of any local learning method in experiments with $N\in\{50,75,100\}$, being outperformed only by backpropagation. With as few as $N=10$ training samples, Forward Projection fitted discriminative models for the classification of OCT (test AUC: $84.5\pm 5.8$) and CXR (test AUC: $76.6\pm 5.8$). CXR and OCT datasets highlighted two distinct vulnerabilities of backpropagation training in few-shot conditions. In the $N=5$ setting on OCT, backpropagation overfitted the training samples (Train AUC: $86.8\pm 13.2$; Test AUC: $61.8\pm 17.4$), as models integrated noise into decision functions. On the other hand, backpropagation failed to achieve adequate model fitting in the $N=10$ setting on CXR (Train AUC: $69.0\pm 8.1$; Test AUC: $59.7\pm 6.2$). Random features performed comparably to backpropagation in tasks with fewest training samples, but performance improved only slightly with larger sample sizes. Random Features models could not overfit within convolutional layers, as these contained no free parameters. However, Random Features had limited capacity to learn structural features in larger training samples. The improvement that Forward Projection provided over Random Features is therefore attributable to structural feature learning within hidden convolutional layers. Label Projection and Noisy Label Projection failed to converge in few-shot training (training AUC $\approx 0.50$). Neither Local Supervision nor Forward-Forward training matched the baseline generalisability of Random Features in OCT and CXR tasks - however Local Supervision achieved comparable test discrimination to Forward Projection in the CIFAR task when $N\geq 75$. Predictive coding and Difference Target Propagation were yielded uninformative models in few-shot learning tasks (Figure 7). Few-shot performance is tabulated in Supplemental Table 4.
@@
Figure 4: Visualisation of layer explanations over space in 2D-convolutional neural networks trained by Forward Projection to detect choroid neovascularization (CNV) in the OCT task. Ensemble average of five models shown. Patients A-C, (diagnosed CNV) and patient D (no disease) were extracted from test data. Explanations were extracted from the second, fourth and sixth convolutional layers ($\hat{y}_{2},\hat{y}_{4},\hat{y}_{6}$), using (3). CNV heat-maps demonstrate high values (red) over CNV features, including retinal/sub-retinal fluid (Patients A-C) and hard exudates (Patient B), and fibrosis (Patient C), with low values (blue) over healthy retina (Patient D).
@@
For backpropagation, $\mathcal{O}(m^{2})$ memory is required for storing layer weights, $\mathcal{O}(m^{2})$ for accumulated gradients, and $\mathcal{O}(m)$ for activations. Here, the computation requirements scale as $\mathcal{O}(N_{e}m^{2})$ MACs each for the forward pass, backward pass, and weight update calculations. Hence, in the typical setting where $m>>m_{L}$, we note that Forward Projection and backpropagation have similar memory requirements, scaling as $\mathcal{O}(m^{2})$. Notably, compute scales as $\mathcal{O}(m^{3})$ for FP versus $\mathcal{O}(N_{e}m^{2})$ for BP. For example, training a dense hidden layer with $1000$ inputs and $1000$ outputs on the FMNIST dataset ($N=60,000,m_{L}=10$) by Forward Projection requires a total of $1.2\times 10^{11}$ MAC operations for the forward pass and $1.2\times 10^{11}$ operations for the weight update. Training the same layer via 100 epochs of backpropagation requires $6.0\times 10^{12}$ MAC operations for forward passes and $1.2\times 10^{13}$ operations for weight updates. Forward projection requires only a single training epoch, leading to a proportional speed up in the overall computation time. Note that Local Supervision and Forward-Forward algorithms have similar computational and memory requirements as backpropagation, as illustrated in Table 2. We also present the wall-clock time for full training of an MLP with $3\times 1000$ hidden neurons on FMNIST, demonstrating a $66\times$ speedup with Forward Projection in this example. Further timing results are provided in Table5 in SI. Computations were run using the Google Colab service with an NVIDIA T4 graphics processing unit.
@@
Backprop-agation $N_{e}$ $N_{e}Nm^{2}$ $2N_{e}Nm^{2}$ $m^{2}$ $m^{2}+2m$ $22.43\pm 4.8$ $11.6\pm 2.7$
Local Supervision $N_{e}$ $N_{e}N(m^{2}+mm_{L})$ $N_{e}N(m^{2}+mm_{L})$ $m^{2}+mm_{L}$ $m^{2}+2m$ $46.9\pm 12.5$ $10.8\pm 3.56$
Forward-Forward $N_{e}$ $2N_{e}Nm^{2}$ $N_{e}Nm^{2}$ $m^{2}$ $m^{2}$ $172\pm 56$ $40.2\pm 13.8$
Predictive Coding $N_{e}$ $N_{e}Nm^{2}$ $4N_{e}Nm^{2}$ $2m^{2}$ $2m^{2}+2m$ $150\pm 61$ $25.4\pm 10.2$
@@
Table 2: Complexities of each training method in an MLP with $m$ inputs, one output and $L$ hidden layers of dimension $m$. NN: training sample size. $N_{e}$: epochs, all parameters except activations are per layer.
@@ Discussion
The closed-form Forward Projection fit is computable in a single pass over the data for each layer, presenting an opportunity to expedite training and reduce environmental footprint. The efficiency of FP training is attributable to a substantially different operational sequence to that employed in SGD-based approaches. Firstly, FP training collects a gram matrix of pre-synaptic activity over a single epoch, which is subsequently inverted during a one-step weight matrix computation. Secondly, FP completes fitting for each layer before initialising successive layers. In contrast, SGD-based approaches fit all the layers in the network iteratively, using feedback from neuronal outputs and downstream layers. FP training requires no retrograde communication between neural output activations and pre-synaptic connections, enabling direct training on hardware with unidirectional synaptic and neuronal communication. This differs from current iterative local learning methods, which require backward communication from neuronal outputs to optimise pre-synaptic parameters (see section A.1 in SI). Although FP addresses the feedback-free constraint of biological learning systems, the biological plausibility of one-step learning through matrix inversion is uncertain. Forward Projection training may be applicable as a pre-training step to reduce the number of training epochs required for backpropagation.
In conclusion, FP is an efficient approach for neural network optimisation, employing techniques from randomised projective embedding and linear regression to fit weight matrices in one epoch using a single-step solution. Interpretability of hidden neurons in FP-trained models may be employed to improve the explainability of neural network predictions.
@@ Methods
The FMNIST dataset was modelled using an MLP with $3\times 1000$ hidden ReLU-activated neurons. Sequential datasets (PTBXL-MI and Promoters) were modelled by a 1d-CNN architecture of four convolutional blocks. Each convolutional block included two convolutional layers with kernel dimension $3$, and strides of 1 and 2 respectively. Convolutional layers in the $l$-th block had $32\times 2^{l-1}$ filters. Convolutional implementation is detailed in Supplementary section A.2. Convolutional outputs were aggregated by global average pooling in the penultimate layer. For gradient-descent-based learning algorithms, batch normalisation layers were included between convolutional blocks. CIFAR modelling in Table 1 used a vision transformer architecture [38] operating on image patches of dimension $4\times 4$, with a sequential stack of four multi-headed attention layers, each having 8 heads, embedding dimension 64, and MLP dimension 64. Few-shot image modelling tasks used a 2d-CNN architecture with kernel dimension $3\times 3$ and either $16\times 2^{l-1}$ filters in the $l$-th convolutional block (CXR and OCT) or $32\times 2^{l-1}$ filters in the $l$-th convolutional block (CIFAR). The $\operatorname{sign}$ function was employed to generate target activations for Forward Projection models, such that

### v3

@@ Abstract
State-of-the-art methods for backpropagation-free learning employ local error feedback to direct iterative optimisation via gradient descent. In this study, we examine the more restrictive setting where retrograde communication from neuronal outputs is unavailable for pre-synaptic weight optimisation. To address this challenge, we propose Forward Projection (FP). This randomised closed-form training method requires only a single forward pass over the entire dataset for model fitting, without retrograde communication. Our method generates target values for pre-activation membrane potentials at each layer through randomised nonlinear projections of pre-synaptic inputs and the labels, thereby encoding information from both sources. Local loss functions are optimised over pre-synaptic inputs using closed-form regression, without feedback from neuronal outputs or downstream layers. Interpretability is a key advantage of FP training; membrane potentials of hidden neurons in FP-trained networks encode information which are interpretable layer-wise as label predictions. We demonstrate the effectiveness of FP across four biomedical datasets, comparing it with backpropagation and local learning techniques such as Forward-Forward training and Local Supervision in multi-layer perceptron and convolutional architectures. In some few-shot learning tasks, FP yielded more generalisable models than those optimised via backpropagation. In large-sample tasks, FP-based models achieve generalisation comparable to gradient descent-based local learning methods while requiring only a single forward propagation step, achieving significant speed up for training. Interpretation functions defined on local neuronal activity in FP-based models successfully identified clinically salient features for diagnosis in two distinct datasets – identification of key signatures of myocardial infarction in time-series data from electrocardiogram sequences and detection of regions of choroid neovascularisation in optical coherence tomography images.
@@ Results
Forward Projection was compared to other local learning methods, including random features (RF) [23], Local Supervision (LS) [9], and Forward-Forward (FF) [15] in multiple tasks using equivalent model architectures (Table 1). The Fashion MNIST (FMNIST) image classification task was modelled on a multi-layer perceptron (MLP). Forward projection achieved higher test accuracy than other local learning methods in this task, approaching the performance of the backpropagation reference standard. Two large-scale biomedical sequence modelling tasks were also evaluated with one-dimensional convolutional neural network (1D-CNN) architectures. The Promoters task required the identification of human non-TATA promoters, a class of gene promoter regions which increase transcription of DNA sequences [26, 27]. Forward projection yielded higher test performance than all other local learning methods in this task. The PTBXL-MI task [28] required diagnosis of myocardial infarction (MI), a heart condition commonly known as “heart attack”, from 12-channel electrocardiograph (ECG) recordings. FP and LS performed comparably in this task. FP also outperformed all other methods in the optimisation of models with more complex neuronal activations such as modulo and polynomial activation functions (see remark A.13 in SI). Backpropagation performance benchmarks in PTBXL-MI and Promoters tasks were comparable to previous studies [29, 27]. For the CIFAR10 classification task without data augmentation, two-dimensional convolutional neural networks (2D-CNNs) trained by FP underfitted, but still outperformed all other local learning methods, whereas standard backpropagation overfitted the model. FP was implemented in a transformer architecture to discriminate the first two classes of the CIFAR10 dataset (“CIFAR2”), significantly outperforming random features.
@@
CIFAR10 2D-CNN AUC 85.1 ± 0.2 81.4 ± 0.5 87.7 ± 0.5 57.9 ± 1.7 74.5 ± 0.3 64.5 ± 2.3 89.8 ± 0.7
Acc 48.8 ± 0.3 40.6 ± 1.0 48.4 ± 1.1 13.1 ± 1.9 28.5 ± 0.9 15.3 ± 3.0 54.8 ± 1.5
Table 1: Test performance of various learning methods across different datasets. The FMNIST dataset was modelled with a multi-layer perception; Promoters and PTBXL-MI datasets were modelled using 1D convolutional neural networks (1D-CNNs); CIFAR2 (first two CIFAR10 classes only) was modelled using vision transformers and CIFAR10 was modelled with 2D-CNNs. BP: bacpropagation; DTP: Difference Target Propagation; $F$: Forward Forward; FP: Forward Projection; LS: Local Supervision; PC: Predictive Coding; RF: random features
@@
Figure 2: A: Comparison of feedback-free fitting methods on FMNIST. MLP architectures had 1000 neurons in the first and second layers and 100, 200, 400 or 800 neurons in the final hidden layer. B-D: Test performance of few-shot trained 2D-CNN models. Mean test AUC is reported over 50 few-shot training experiments. B: Chest X-ray (CXR) task. C: Optical Coherence Tomography (OCT) task. D: CIFAR2 task in which models were required to classify the first two classes (aeroplane and automobile). Mean test AUC is reported over 50 few-shot training experiments. Models were fitted with $N\in\{5,10,15,20,30,40,50\}$ training samples from each class in CXR and OCT tasks and $N\in\{25,50,75,100\}$ samples per class for the CIFAR2 task. Predictive Coding and Difference Target Propagation are plotted separately in Supplementary Figure 7. BP: backpropagation; $F$: Forward-Forward; FP: Forward Projection; LS: Local Supervision; RF: random features.
@@
“Few-shot” learning is a constrained learning scenario in which the number of data samples available for training is small. High-dimensional data, such as images, pose a challenge for few-shot learning, as many spurious features may exist. Thus, successful few-shot training methods must select generalisable features in the presence of these confounders. Few-shot learning was assessed in image classification tasks using a 2D-CNN architecture. The optical coherence tomography task (OCT) [30] required that the model discriminate images of healthy retinas from those with choroid neovascularization, a pathology that affects the eye and manifests mainly as abnormal growth of blood vessels behind the retina. The paediatric chest X-ray (CXR) task [30] required the model to discriminate between images of viral pneumonia, bacterial pneumonia, or healthy controls. The CIFAR2 task required the model to discriminate between the first two CIFAR10 classes (aeroplane and automobile). Few-shot training datasets were generated by subsampling, with $N\in\{5,10,15,20,30,40,50\}$ training examples from each class for OCT and CXR tasks, and $N\in\{25,50,75,100\}$ training examples per class for the CIFAR2 task. Data augmentation was not employed. Model generalisability was assessed on all test data (CXR: $N_{\text{test}}=431$; OCT: $N_{\text{test}}=327$, CIFAR2: $N_{\text{test}}=2000$). Forward Projection-trained models demonstrated the greatest few-shot generalisability in CXR (Figure 2-B), OCT (Figure 2-C), outperforming all other methods, including backpropagation, in all the tasks with $N\leq 40$ training samples. In the CIFAR few-shot task, Forward Projection achieved the best performance of any local learning method in experiments with $N\in\{50,75,100\}$, being outperformed only by backpropagation. With as few as $N=10$ training samples, Forward Projection fitted discriminative models for the classification of OCT (test AUC: $84.5\pm 5.8$) and CXR (test AUC: $76.6\pm 5.8$). CXR and OCT datasets highlighted two distinct vulnerabilities of backpropagation training in few-shot conditions. In the $N=5$ setting on OCT, backpropagation overfitted the training samples (Train AUC: $86.8\pm 13.2$; Test AUC: $61.8\pm 17.4$), as models integrated noise into decision functions. On the other hand, backpropagation failed to achieve adequate model fitting in the $N=10$ setting on CXR (Train AUC: $69.0\pm 8.1$; Test AUC: $59.7\pm 6.2$). Random features performed comparably to backpropagation in tasks with fewest training samples, but performance improved only slightly with larger sample sizes. Random Features models could not overfit within convolutional layers, as these contained no free parameters. However, Random Features had limited capacity to learn structural features in larger training samples. The improvement that Forward Projection provided over Random Features is therefore attributable to structural feature learning within hidden convolutional layers. Label Projection and Noisy Label Projection failed to converge in few-shot training (training AUC $\approx 0.50$). Neither Local Supervision nor Forward-Forward training matched the baseline generalisability of Random Features in OCT and CXR tasks - however Local Supervision achieved comparable test discrimination to Forward Projection in the CIFAR2 task when $N\geq 75$. Predictive coding and Difference Target Propagation were yielded uninformative models in few-shot learning tasks (Figure 7). Few-shot performance is tabulated in Supplemental Table 4.
@@
Figure 4: Visualisation of layer explanations over space in 2D-CNNs trained by Forward Projection to detect choroid neovascularization (CNV) in the OCT task. Ensemble average of five models shown. Patients A-C, (diagnosed CNV) and patient D (no disease) were extracted from test data. Explanations were extracted from the second, fourth and sixth convolutional layers ($\hat{y}_{2},\hat{y}_{4},\hat{y}_{6}$), using (3). CNV heat-maps demonstrate high values (red) over CNV features, including retinal/sub-retinal fluid (Patients A-C) and hard exudates (Patient B), and fibrosis (Patient C), with low values (blue) over healthy retina (Patient D).
@@
For backpropagation, $\mathcal{O}(m^{2})$ memory is required for storing layer weights, $\mathcal{O}(m^{2})$ for accumulated gradients, and $\mathcal{O}(m)$ for activations. Here, the computation requirements scale as $\mathcal{O}(N_{e}m^{2})$ MACs each for the forward pass, backward pass, and weight update calculations. Hence, in the typical setting where $m>>m_{L}$, we note that Forward Projection and backpropagation have similar memory requirements, scaling as $\mathcal{O}(m^{2})$. Notably, compute scales as $\mathcal{O}(m^{3})$ for FP versus $\mathcal{O}(N_{e}m^{2})$ for BP. For example, training a dense hidden layer with $1000$ inputs and $1000$ outputs on the FMNIST dataset ($N=60,000,m_{L}=10$) by Forward Projection requires a total of $1.2\times 10^{11}$ MAC operations for the forward pass and $1.2\times 10^{11}$ operations for the weight update. Training the same layer via 100 epochs of backpropagation requires $6.0\times 10^{12}$ MAC operations for forward passes and $1.8\times 10^{13}$ operations for weight updates. Forward projection requires only a single training epoch, leading to a proportional speed up in the overall computation time. Note that Local Supervision and Forward-Forward algorithms have similar computational and memory requirements as backpropagation, as illustrated in Table 2. We also present the wall-clock time for full training of an MLP with $3\times 1000$ hidden neurons on FMNIST, demonstrating a $66\times$ speedup with Forward Projection in this example. Further timing results are provided in Table5 in SI. Computations were run using the Google Colab service with an NVIDIA T4 graphics processing unit. Training complexities of other methods are discussed in remark A.12.
@@
Backprop-agation $N_{e}$ $N_{e}Nm^{2}$ $3N_{e}Nm^{2}$ $m^{2}$ $m^{2}+2m$ $22.43\pm 4.8$ $11.6\pm 2.7$
Local Supervision $N_{e}$ $N_{e}N(m^{2}+mm_{L})$ $2N_{e}N(m^{2}+mm_{L})$ $m^{2}+mm_{L}$ $m^{2}+2m$ $46.9\pm 12.5$ $10.8\pm 3.56$
Forward-Forward $N_{e}$ $2N_{e}Nm^{2}$ $3N_{e}Nm^{2}$ $m^{2}$ $m^{2}+2m$ $172\pm 56$ $40.2\pm 13.8$
Predictive Coding $N_{e}$ $N_{e}Nm^{2}$ $3N_{e}Nm^{2}$ $2m^{2}$ $2m^{2}+2m$ $150\pm 61$ $25.4\pm 10.2$
@@
Table 2: Training complexity for a single hidden layer with $m$ inputs, $m$ outputs, given a label with dimension $m_{L}$. NN: training sample size. $N_{e}$: epochs, all parameters except activations are per layer. For simplicity, we assume batch size $B=1$ and that predictive coding trains for $k=1$ iteration. Complexity is detailed in remark A.12 in SI.
@@ Discussion
The closed-form Forward Projection fit is computable in a single pass over the data for each layer, presenting an opportunity to expedite training and reduce environmental footprint. The efficiency of FP training is attributable to a substantially different operational sequence to that employed in SGD-based approaches. Firstly, FP training collects a gram matrix of pre-synaptic activity over a single epoch, which is subsequently inverted during a one-step weight matrix computation. Secondly, FP completes fitting for each layer before initialising successive layers. In contrast, SGD-based approaches fit all the layers in the network iteratively, using feedback from neuronal outputs and downstream layers. FP training requires no retrograde communication between neural output activations and pre-synaptic connections, enabling direct training on hardware with unidirectional synaptic and neuronal communication. This differs from current iterative local learning methods, which require backward communication from neuronal outputs to optimise pre-synaptic parameters (see remark A.1 in SI). Although FP addresses the feedback-free constraint inherent in biological learning systems, the biological plausibility of one-step learning via matrix inversion remains uncertain. Iterative fitting via gradient descent in sequential, batch-based learning offers a practical and scalable approach for aligning weights with locally generated targets, as defined in (4) (see remark A.10 in SI). Although closed-form optimisation remains a challenge for recurrent architectures, future research will explore integrating iterative training with random temporal convolutions to generalise the FP framework to dynamic models and expand its applicability to temporal learning tasks. This backpropagation-free strategy holds promise for biologically inspired computing systems, such as spiking neural networks, where non-differentiable activation functions preclude standard backpropagation [36, 37, 38]. FP training may be applicable as a pre-training step to reduce the number of training epochs required for backpropagation. In conclusion, FP is an efficient approach for neural network optimisation, employing techniques from randomised projective embedding and linear regression to fit weight matrices in one epoch using a single-step solution. Interpretability of hidden neurons in FP-trained models may be employed to improve the explainability of neural network predictions.
@@ Methods
The FMNIST dataset was modelled using an MLP with $3\times 1000$ hidden ReLU-activated neurons. Sequential datasets (PTBXL-MI and Promoters) were modelled by a 1D-CNN architecture of four convolutional blocks. Each convolutional block included two convolutional layers with kernel dimension $3$, and strides of 1 and 2 respectively. Convolutional layers in the $l$-th block had $32\times 2^{l-1}$ filters. Convolutional implementation is detailed in Supplementary section A.2. Convolutional outputs were aggregated by global average pooling in the penultimate layer. For gradient-descent-based learning algorithms, batch normalisation layers were included between convolutional blocks. CIFAR2 modelling in Table 1 used a vision transformer architecture [41] operating on image patches of dimension $4\times 4$, with a sequential stack of four multi-headed attention layers, each having 8 heads, embedding dimension 64, and MLP dimension 64. 2D-CNN architectures had kernel dimension $3\times 3$and the $l$-th convolutional block had $16\times 2^{l-1}$ filters (CXR and OCT) $32\times 2^{l-1}$ filters (CIFAR2) or $64\times 2^{l-1}$ filters (CIFAR10). The $\operatorname{sign}$ function was employed to generate target activations for Forward Projection models, such that